OpenStudy (pink.1d):

Choose a row and column and compare P(A | B) with P(B | A). Explain what each probability means in the context of the situation and data you collected.

OpenStudy (pink.1d):

This is my graph

OpenStudy (pink.1d):

table***

OpenStudy (pink.1d):

Use social media regularly Don't use social media Younger than 18 16 4 Older than 18 11 9

OpenStudy (pink.1d):

@perl

OpenStudy (pink.1d):

also: Compare P(A∩B) with P(A∪B), and explain what each probability means in the context of the situation and data you collected. SAME TABLE posted above

OpenStudy (pink.1d):

@dan815 @iambatman @triciaal

OpenStudy (perl):

lets start with the first question

OpenStudy (pink.1d):

okay:)

OpenStudy (perl):

first find the totals http://prntscr.com/7cj8yl

OpenStudy (perl):

Let A = Use social media regularly Let B = Younger than 18 P( A | B) = 16/20 P(B|A) = 16/27

OpenStudy (pink.1d):

i understand up to now

OpenStudy (pink.1d):

i dont get what they mean when it says ' Explain what each probability means in the context of the situation and data you collected.' i dont know how to answer that. i know how to find the answer but i m not sure how to answer that

OpenStudy (perl):

P(A | B) means probability use social media regularly given that we only look at people who are younger than 18 http://prntscr.com/7cja04 Or another way to say it, of people who are younger than 18 what is probability that randomly selecting one uses social media

OpenStudy (perl):

so we only look at that top row

OpenStudy (pink.1d):

so the top row and 1s2 colum?

OpenStudy (pink.1d):

1st*

OpenStudy (pink.1d):

younger than 18 and use social media

OpenStudy (pink.1d):

@perl @perl @perl @perl dont leave

OpenStudy (pink.1d):

YAY ur back! :)

OpenStudy (perl):

right, top row and first colmun is P(A|B)

OpenStudy (perl):

P(A|B) corresponds to top row and first colmun also we can express it as P(A|B) =P(A∩B) / P(B)

OpenStudy (pink.1d):

what does p(b line a) mean. it means

OpenStudy (perl):

P(B|A) says to look in only first column, first row

OpenStudy (pink.1d):

so what exactly does that mean?

OpenStudy (pink.1d):

im confused

OpenStudy (pink.1d):

does it mean it all out of total

OpenStudy (perl):

If we define http://prntscr.com/7cjcnn P(B|A) says to look in only first column A P(A|B) says to look only in first row B

OpenStudy (pink.1d):

can you look over the answer I wrote?

OpenStudy (perl):

what did you write

OpenStudy (pink.1d):

P(B | A) means probability of those under 18 who use social media regularly out of the total people that use social media regularly

OpenStudy (pink.1d):

P(A | B) : This probability means means probability of those under 18 who use social media regularly out of all the people under 18

OpenStudy (perl):

do you agre that I defined A = "those who use social media regularly"

OpenStudy (pink.1d):

am i right?

OpenStudy (perl):

agree*

OpenStudy (pink.1d):

or wrong

OpenStudy (perl):

its kind of hard to say, its a bit confusing what you wrote for the second one

OpenStudy (pink.1d):

P(A | B) : This means probability of those under 18 who use social media regularly out of all the people under 18

OpenStudy (pink.1d):

i edited that

OpenStudy (pink.1d):

is the first one right?

OpenStudy (pink.1d):

@perl

OpenStudy (perl):

P(A|B): probability that a person who is younger than 18 uses social media regularly.

OpenStudy (perl):

P(B|A): Probability that a person who uses social media regularly is younger than 18.

OpenStudy (perl):

http://prntscr.com/7cjfig

OpenStudy (pink.1d):

@perl can you help me with the second one please?

OpenStudy (pink.1d):

NEVERMIND

OpenStudy (pink.1d):

THANKS FOR ALL THE HELP